If the sum of a number and 5 is doubled, the result is 4 less than the number. find the number.I tried this and it didn't work out correctly.
2(x+5)=2x-4
2x+10=2x-4
2x+10-2x=2x-4-2x
now ive got
10=-4
10-4=6
this isn't the correct answer where did I screw up or forget something?
I would appreciate your help

Answer by lwsshak3(11628)

If the sum of a number and 5 is doubled, the result is 4 less than the number. find the number.I tried this and it didn't work out correctly.
**
let x=number
2(x+5)=x-4
2x+10=x-4
x=-14